# Concept – An Evaluation Protocol on Conversation Recommender Systems with System- and User-centric Factors

### v1

@@ 5 Conclusion
We regard CRS as a social issue, instead of just a technical problem, considering its major goal of persuading users to accept the recommendations. Hence, we revisit current evaluation protocols and pioneer works to provide a comprehensive survey on understanding the impact of CRS’s characteristics in terms of both pros and cons. We then come up with a new evaluation protocol– Concept, which considers both system- and user-centric factors for addressing personalized user needs in particular, with the hope of enhancing user experience in CRS. Such factors include three key characteristics, which are further divided into six primary abilities. In addition, we suggest that even enhanced by "omnipotent" LLMs, CRS lacks practical usability. We pinpoint several significant limitations of current CRS models, including e.g. the incapacity to express itself sincerely, a lack of identity-awareness, low reliability towards contextual nuance, and poor coordination for diverse users. Through Concept, we aim for providing CRS developers and practitioners with a guidance to draw meaning insights from our inclusive and fine-grained evaluation directions.
6 Limitations
@@
The use of LLM-based user simulators and evaluators is a double-edged sword. It is a labor-intensive and effective approach, but may suffer from weak robustness by its nature. Although we have adopted some strategies to improve the robustness as much as possible following previous works, the robustness may still be hampered due to the potential uncertainty of prompt engineering. Additionally, another aspect that could improve the Concept is to generate more conversation data, run the LLM-based evaluation many times with different seeds, and report more statistically significant results. However, budgeting is always a factor to consider. In this case, it is important to propose a user simulator and evaluator based on an open-source small model that has similar capabilities to ChatGPT. Finally, our current work does not evaluate the attribute-based CRS Lei et al. (2020), as this type of research often ignores the ability to engage in smooth conversations and instead focuses solely on accurately providing recommendations to users within minimal conversation turns. In this case, evaluating attribute-based CRS models seems unfair. We highlight the importance of combining attribute-based and dialog-based CRS studies to create a more holistic CRS, taking into account its practical usability.
Appendix A Implementation Details
@@
We conduct all our experiments using a single Nvidia RTX A6000, and we implement our codes in PyTorch. The core of our code framework is built upon open-source code from the latest research on CRS [^10] https://github.com/RUCAIBox/iEvaLM-CRS Wang et al. (2023d). In order to guarantee replicability, we have established fixed values for the Temperature and Seed parameters of ChatGPT (i.e., GPT-3.5-16K-turbo), setting the Temperature to 0 and the Seed to 42.
A.1 Implementation of CRS Models
@@
We evaluated all open-source CRS models in its code using their checkpoints, including the KBRD Chen et al. (2019), BARCOR Wang et al. (2022a), UNICRS Wang et al. (2022b), and CHATGPT-based model Wang et al. (2023d) which represents the current SOTA, incorporating text-embedding-ada-002 Neelakantan et al. (2022) for modeling the semantic embeddings.
@@
  • KBRD Chen et al. (2019) bridges the recommendation module and the Transformer-based conversation module via knowledge propagation.
@@
  • BARCOR Wang et al. (2022a) is a unified framework based on BART Lewis et al. (2020), which implements the recommendation and response generation tasks in a single model.
@@
  • UNICRS Wang et al. (2022b) is a unified framework based on DialoGPT Zhang et al. (2020), with a semantic fusion module to enhance the semantic association between conversation history and knowledge graphs.
@@
  • CHATCRS Wang et al. (2023d) is the SOTA CRS model, which incorporates ChatGPT for the conversation module and the text-embedding-ada-002 Neelakantan et al. (2022) to enhance the recommendation module.
@@
See the prompts in the original paper for the implementation of CHATGPT-based model. Note that we feed the top 5 items from the recommendation module into the prompts of ChatGPT for re-ranking and generating responses. Also, the UNICRS utilizes DialoGPT-small as the backbone, while the BARCOR utilizes BART-base with a 2-layer encoder and decoder, following Wang et al. (2023d).
@@ A.2.1 Evaluation Process
Afterward, Concept utilizes both the LLM-based evaluator and computational metrics to assess the abilities of CRS, as outlined in Table 1. Note that the evaluator is utilized when corresponding computational metrics are not available. Specifically, we employ the evaluator to assess the cooperation ability, social awareness, and persuasiveness of recommendation explanations of CRS. This involves prompting the evaluator with fine-grained, ability-specific scoring rubrics Saha et al. (2023b), which are generated by the LLM and then refined by humans. For more details, refer to Appendix A.2.3 and Appendix A.2.4.
A.2.2 LLM-based User Simulator
Our user simulator has unique personas and preferences. Specifically, the personas are generated by prompting ChatGPT in a zero-shot manner, following (Wang et al., 2023a), while the preferences are defined using attributes from two benchmark datasets, Redial and OpendialKG.
Personas. We generate a persona list as the starting point by prompting ChatGPT in a zero-shot manner. Inspired by Wang et al. (2023a), we first instruct the ChatGPT to return 20 distinct personas and their corresponding descriptions, such as curiosity, and then filter out duplicates. We utilize the generated persona list, so it is highly likely that ChatGPT is familiar with these personas and would contain knowledge and information about them. As a result, we obtain 12 unique personas, simulating diverse sentiments may encountered when using a conversational recommendation system (namely, Anticipation, Boredom, Fusion, Curiosity, Delight, Disappointment, Exceptions, Frustration, Independence, Surprise, Trust, Satisfaction), and each type of user may belong to different age groups (namely, Adults, Children, Senior, and Teens). We assigned different attribute groups as movie preferences for each user type. These users with varying emotions, age groups, and movie preferences then interacted with the CRS. If the attribute group of the movie recommended by the CRS is equal to or includes the user’s target attribute group, it is considered a correct recommendation. For the Redial dataset, each user type corresponds to 76 conversation data with one CRS, resulting in a total of 912 conversation data across all user types and one CRS. Similarly, for the OpendialKG dataset, each user type generates 64 conversation data with each CRS, resulting in a total of 768 conversation data across all user types and one CRS. Since we tested 4 different CRSs, we collected a total of 6720 conversation data. Refer to Table 17 for details.
@@ A.2.2 LLM-based User Simulator
Preferences. The current evaluation method for CRS is not reflective of real-world situations, as it assumes that every user knows the target item Wang et al. (2023d). To tackle this issue, our user simulators only have access to their preferences. The user would only choose to accept the recommendation when the CRS provides explanations for the recommendations and makes the users feel that the recommended items match their preferences. To achieve this, the preferences are defined using attributes from two benchmark datasets, Redial Redial Li et al. (2018) and OpendialKG Moon et al. (2019). For each dataset, every movie has a feature group made up of one or more attributes. In the case of the Redial dataset, we conducted experiments using feature groups containing 3 attributes and excluded less common groups. Ultimately, we retained the 19 most prevalent attribute groups for the study, with each group corresponding to at least 50 different movies. For the OpendialKG dataset, the issue of less common attributes is more pronounced. Initially, we selected the most prevalent attributes (each corresponding to at least 100 movies) and then kept the 16 most common attribute groups for experimentation.
Simulation via Prompts. We prompt the simulator with different personas using the persona descriptions generated by ChatGPT (cf. Table 7.). Concept incorporates the Theory of Mind into our simulator to emulate human social cognition Fischer (2023). This is achieved by prompting the simulator to first assess its current mental state before generating responses, encouraging reflection on its predefined personality traits and social interactions. Refer to Appendix F for prompts. Additionally, there is a maximum of 10 turns allowed in the conversation, and it will only conclude when the simulated user accepts the recommendation. To prevent the simulated users from directly stating the same attribute preferences as the pre-defined values (i.e., the attribute group) when asked about their preferences by the CRS, we include adjectives before each attribute during the simulation, and we allow the user to describe his/her preference using their own words. This approach is reasonable because in real scenarios, users may not use the exact words as the pre-defined attribute values during the conversation. We have summarized the adjusted values for each attribute in Tables 6, achieved by prompting ChatGPT.
A.2.3 LLM-based Evaluator
@@
Based on the findings of previous studies Ye et al. (2023); Wang et al. (2023c, d), we utilize an instance-wise evaluator to conduct detailed assessments. Expanding on earlier research Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023d), we task the evaluator with using fine-grained scoring rubrics to mitigate scoring bias (see Section D for our reliability analysis). The evaluator assigns scores ranging from 1 to 5 based on ability-specific score rubrics, each accompanied by a corresponding description, using the conversation data. For the generation of fine-grained scoring rubrics, we follow the approach of previous works and employ ChatGPT to produce a set of evaluation criteria Saha et al. (2023a). This set of criteria serves as a starting point for human refinement Li et al. (2023b). Prior to assigning a score, we require the evaluator to provide a rationale, drawing inspiration from the effectiveness of CoT prompting Wei et al. (2022); Ye et al. (2023). For implementation details on the evaluator, refer to Appendix F.
@@
Table 7: Persona description generation. We start by converting each combination of personality and age into a single sentence using a template, and then we use ChatGPT to rephrase it into a natural-sounding sentence. Refer to Appendix F for Prompts.
A.2.4 Computational Metrics
@@
  • Quality. We consider computational metrics for automatic evaluation following previous works Wang et al. (2023d); Zhang et al. (2023); Yu et al. (2023), including $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate ($AR$), and average turns ($AT$) required to reach the successful recommendation. Note that we determine user acceptance of the recommendation by checking for the presence of the special token ’[END]’ in the user’s response.
@@
CRS-Que Jin et al. (2023)
CRS-UX Jin et al. (2021)
US Siro et al. (2023)
INSPIRED Hayati et al. (2020)
iEval Wang et al. (2023d)
@@
Appendix B More Related Works
@@
CRS Evaluation. The practical effectiveness of CRS has always been questioned Jannach and Manzoor (2020); Jannach et al. (2021). Addressing this concern necessitates an inclusive evaluation protocol to assess the strengths, weaknesses, and potential risks of CRS. The success of CRS in practice relies primarily on social characteristics, not just technical ones. However, most evaluation protocols primarily focus on superficial and limited evaluation aspects (in isolation) from system-centric perspectives, such as lexical diversity and perplexity of responses Ghazvininejad et al. (2018); Chen et al. (2019), conversational fluency, relevance, informativeness Wang et al. (2022b, a), recommendation effectiveness and efficiency Wang et al. (2023d); Jin et al. (2019); Wärnestål (2005), and rapport building Hayati et al. (2020). Since they fail to reveal the inadequacy of CRS from user-centric perspectives, these fragmented evaluation protocols are underdeveloped to constitute a standard benchmark and conduct an inclusive evaluation. Recent efforts have attempted to systematically summarize and evaluate the characteristics of CRS, but they have mainly centered on establishing evaluation protocols through person-to-person conversation analysis and questionnaire interviews, lacking quantitative observations on existing CRS models Jin et al. (2023, 2021); Jannach (2022); Siro et al. (2023) and failing to systematically evaluate system-centric characteristics. Refer to Table 8 for better understanding our differences. In summary, existing protocols are considered too superficial and inadequate to uncover the shortcomings of CRS Jannach et al. (2021); Jannach (2023). To this end, our Concept brings together system- and user-centric perspectives, enabling a more profound and inclusive evaluation protocol and ultimately contributing to the development of more practical CRS Li et al. (2023a).
@@
LLM as User Simulator. Manual evaluation often requires the user to constantly interact with CRS, which are labor-expensive and generally only feasible in an industry lab Huang et al. (2023); Gao et al. (2021). To this end, LLM-based user simulators continue to be a rapidly evolving research area due to their labor-intensive nature and effectiveness Kocmi and Federmann (2023); Terragni et al. (2023); Tang et al. (2023); Cheng et al. (2023); Afzali et al. (2023). A recent study on CRS also features interactive evaluation with LLM-based user simulation Wang et al. (2023d), experimentally demonstrating its effectiveness and reliability as a reliable alternative to humans. Taking inspiration from these studies, Concept also harnesses an LLM-based user simulator to interact with the CRS. Moreover, we equip our simulator with elaborately designed instructions and the Theory of Mind to mimic human social cognition Fischer (2023). This encourages the simulator to reflect on its own predefined personality traits.
@@
LLM as Evaluator. A LLM-based evaluator could be seen as a combination of a LLM and its prompting strategy Zeng et al. (2023); Cohen et al. (2023); Chan et al. (2023); Wang et al. (2023b); Liu et al. (2023b). Considering its flexibility, the instance-wise evaluator is widely adopted Ye et al. (2023); Wang et al. (2023c); Liu et al. (2023a). Importantly, previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. (2023d). These findings are further supported by recent work on CRS Wang et al. (2023d), resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. Inspired by them, we involve ability-specific scoring rubrics to gain reliable and aligned evaluations. We also encourage the evaluator to generate a rationale before assigning a score, further ensuring reliability Ye et al. (2023).
Appendix C Characteristics Categorization

### v2

@@ 5 Conclusion
We regard CRS as a social issue, instead of just a technical problem, considering its major goal of persuading users to accept the recommendations. Hence, we revisit current evaluation protocols and pioneer works to provide a comprehensive survey on understanding the impact of CRS’s characteristics in terms of both pros and cons. We then come up with a new evaluation protocol– Concept, which considers both system- and user-centric factors for addressing personalized user needs in particular, with the hope of enhancing user experience in CRS. Such factors include three key characteristics, which are further divided into six primary abilities. In addition, we suggest that even enhanced by "omnipotent" LLMs, CRS lacks practical usability. We pinpoint several significant limitations of current CRS models, including the incapacity to express itself sincerely, a lack of identity-awareness, low reliability towards contextual nuance, and poor coordination for diverse users. However, our current work does not evaluate the attribute-based CRS Lei et al. (2020), as this type of research often ignores the ability to engage in smooth conversations and instead focuses solely on accurately providing recommendations to users within minimal conversation turns. In this case, evaluating attribute-based CRS models seems unfair. To sum up, through Concept, we aim for providing CRS developers and practitioners with a guidance to draw meaning insights from our inclusive and fine-grained evaluation directions.
Appendix A Limitation Analysis of Our Work
@@
The use of LLM-based user simulators and evaluators is a double-edged sword. It is a labor-intensive and effective approach, but may suffer from weak robustness by its nature. Although we have adopted some strategies to improve the robustness as much as possible following previous works, the robustness may still be hampered due to the potential uncertainty of prompt engineering. Additionally, another aspect that could improve the Concept is to generate more conversation data, run the LLM-based evaluation many times with different seeds, and report more statistically significant results. However, budgeting is always a factor to consider. In this case, it is important to propose a user simulator and evaluator based on an open-source small model that has similar capabilities to ChatGPT. Finally, our current work does not evaluate the attribute-based CRS Lei et al. [2020], as this type of research often ignores the ability to engage in smooth conversations and instead focuses solely on accurately providing recommendations to users within minimal conversation turns. In this case, evaluating attribute-based CRS models seems unfair. We highlight the importance of combining attribute-based and dialog-based CRS studies to create a more holistic CRS, taking into account its practical usability.
Appendix B Implementation Details
@@
We conduct all our experiments using a single Nvidia RTX A6000, and we implement our codes in PyTorch. The core of our code framework is built upon open-source code from the latest research on CRS [^11] https://github.com/RUCAIBox/iEvaLM-CRS Wang et al. [2023d]. In order to guarantee replicability, we have established fixed values for the Temperature and Seed parameters of ChatGPT (i.e., GPT-3.5-16K-turbo), setting the Temperature to 0 and the Seed to 42.
B.1 Implementation of CRS Models
@@
We evaluated all open-source CRS models in its code using their checkpoints, including the KBRD Chen et al. [2019], BARCOR Wang et al. [2022a], UNICRS Wang et al. [2022b], and CHATGPT-based model Wang et al. [2023d] which represents the current SOTA, incorporating text-embedding-ada-002 Neelakantan et al. [2022] for modeling the semantic embeddings.
@@
  • KBRD Chen et al. [2019] bridges the recommendation module and the Transformer-based conversation module via knowledge propagation.
@@
  • BARCOR Wang et al. [2022a] is a unified framework based on BART Lewis et al. [2020], which implements the recommendation and response generation tasks in a single model.
@@
  • UNICRS Wang et al. [2022b] is a unified framework based on DialoGPT Zhang et al. [2020], with a semantic fusion module to enhance the semantic association between conversation history and knowledge graphs.
@@
  • CHATCRS Wang et al. [2023d] is the SOTA CRS model, which incorporates ChatGPT for the conversation module and the text-embedding-ada-002 Neelakantan et al. [2022] to enhance the recommendation module.
@@
See the prompts in the original paper for the implementation of CHATGPT-based model. Note that we feed the top 5 items from the recommendation module into the prompts of ChatGPT for re-ranking and generating responses. Also, the UNICRS utilizes DialoGPT-small as the backbone, while the BARCOR utilizes BART-base with a 2-layer encoder and decoder, following Wang et al. [2023d].
@@ B.2.1 Evaluation Process
Afterward, Concept utilizes both the LLM-based evaluator and computational metrics to assess the abilities of CRS, as outlined in Table 1. Note that the evaluator is utilized when corresponding computational metrics are not available. Specifically, we employ the evaluator to assess the cooperation ability, social awareness, and persuasiveness of recommendation explanations of CRS. This involves prompting the evaluator with fine-grained, ability-specific scoring rubrics Saha et al. [2023b], which are generated by the LLM and then refined by humans. For more details, refer to Appendix B.2.3 and Appendix B.2.4.
B.2.2 LLM-based User Simulator
Our user simulator has unique personas and preferences. Specifically, the personas are generated by prompting ChatGPT in a zero-shot manner, following [Wang et al., 2023a], while the preferences are defined using attributes from two benchmark datasets, Redial and OpendialKG.
Personas. We generate a persona list as the starting point by prompting ChatGPT in a zero-shot manner. Inspired by Wang et al. [2023a], we first instruct the ChatGPT to return 20 distinct personas and their corresponding descriptions, such as curiosity, and then filter out duplicates. We utilize the generated persona list, so it is highly likely that ChatGPT is familiar with these personas and would contain knowledge and information about them. As a result, we obtain 12 unique personas, simulating diverse sentiments may encountered when using a conversational recommendation system (namely, Anticipation, Boredom, Fusion, Curiosity, Delight, Disappointment, Exceptions, Frustration, Independence, Surprise, Trust, Satisfaction), and each type of user may belong to different age groups (namely, Adults, Children, Senior, and Teens). We assigned different attribute groups as movie preferences for each user type. These users with varying emotions, age groups, and movie preferences then interacted with the CRS. If the attribute group of the movie recommended by the CRS is equal to or includes the user’s target attribute group, it is considered a correct recommendation. For the Redial dataset, each user type corresponds to 76 conversation data with one CRS, resulting in a total of 912 conversation data across all user types and one CRS. Similarly, for the OpendialKG dataset, each user type generates 64 conversation data with each CRS, resulting in a total of 768 conversation data across all user types and one CRS. Since we tested 4 different CRSs, we collected a total of 6720 conversation data. Refer to Table 17 for details.
@@ B.2.2 LLM-based User Simulator
Preferences. The current evaluation method for CRS is not reflective of real-world situations, as it assumes that every user knows the target item Wang et al. [2023d]. To tackle this issue, our user simulators only have access to their preferences. The user would only choose to accept the recommendation when the CRS provides explanations for the recommendations and makes the users feel that the recommended items match their preferences. To achieve this, the preferences are defined using attributes from two benchmark datasets, Redial Redial Li et al. [2018] and OpendialKG Moon et al. [2019]. For each dataset, every movie has a feature group made up of one or more attributes. In the case of the Redial dataset, we conducted experiments using feature groups containing 3 attributes and excluded less common groups. Ultimately, we retained the 19 most prevalent attribute groups for the study, with each group corresponding to at least 50 different movies. For the OpendialKG dataset, the issue of less common attributes is more pronounced. Initially, we selected the most prevalent attributes (each corresponding to at least 100 movies) and then kept the 16 most common attribute groups for experimentation.
Simulation via Prompts. We prompt the simulator with different personas using the persona descriptions generated by ChatGPT (cf. Table 7.). Concept incorporates the Theory of Mind into our simulator to emulate human social cognition Fischer [2023]. This is achieved by prompting the simulator to first assess its current mental state before generating responses, encouraging reflection on its predefined personality traits and social interactions. Refer to Appendix G for prompts. Additionally, there is a maximum of 10 turns allowed in the conversation, and it will only conclude when the simulated user accepts the recommendation. To prevent the simulated users from directly stating the same attribute preferences as the pre-defined values (i.e., the attribute group) when asked about their preferences by the CRS, we include adjectives before each attribute during the simulation, and we allow the user to describe his/her preference using their own words. This approach is reasonable because in real scenarios, users may not use the exact words as the pre-defined attribute values during the conversation. We have summarized the adjusted values for each attribute in Tables 6, achieved by prompting ChatGPT.
B.2.3 LLM-based Evaluator
@@
Based on the findings of previous studies Ye et al. [2023], Wang et al. [2023c, d], we utilize an instance-wise evaluator to conduct detailed assessments. Expanding on earlier research Ye et al. [2023], Wang et al. [2023d], Liu et al. [2023d], we task the evaluator with using fine-grained scoring rubrics to mitigate scoring bias (see Section E for our reliability analysis). The evaluator assigns scores ranging from 1 to 5 based on ability-specific score rubrics, each accompanied by a corresponding description, using the conversation data. For the generation of fine-grained scoring rubrics, we follow the approach of previous works and employ ChatGPT to produce a set of evaluation criteria Saha et al. [2023a]. This set of criteria serves as a starting point for human refinement Li et al. [2023b]. Prior to assigning a score, we require the evaluator to provide a rationale, drawing inspiration from the effectiveness of CoT prompting Wei et al. [2022], Ye et al. [2023]. For implementation details on the evaluator, refer to Appendix G.
@@
Table 7: Persona description generation. We start by converting each combination of personality and age into a single sentence using a template, and then we use ChatGPT to rephrase it into a natural-sounding sentence. Refer to Appendix G for Prompts.
B.2.4 Computational Metrics
@@
  • Quality. We consider computational metrics for automatic evaluation following previous works Wang et al. [2023d], Zhang et al. [2023], Yu et al. [2023], including $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate ($AR$), and average turns ($AT$) required to reach the successful recommendation. Note that we determine user acceptance of the recommendation by checking for the presence of the special token ’[END]’ in the user’s response.
@@
CRS-Que Jin et al. [2023]
CRS-UX Jin et al. [2021]
US Siro et al. [2023]
INSPIRED Hayati et al. [2020]
iEval Wang et al. [2023d]
@@
Appendix C More Detailed Related Works
@@
CRS Evaluation. The practical effectiveness of CRS has always been questioned Jannach and Manzoor [2020], Jannach et al. [2021]. Addressing this concern necessitates an inclusive evaluation protocol to assess the strengths, weaknesses, and potential risks of CRS. The success of CRS in practice relies primarily on social characteristics, not just technical ones. However, most evaluation protocols primarily focus on superficial and limited evaluation aspects (in isolation) from system-centric perspectives, such as lexical diversity and perplexity of responses Ghazvininejad et al. [2018], Chen et al. [2019], conversational fluency, relevance, informativeness Wang et al. [2022b, a], recommendation effectiveness and efficiency Wang et al. [2023d], Jin et al. [2019], Wärnestål [2005], and rapport building Hayati et al. [2020]. Since they fail to reveal the inadequacy of CRS from user-centric perspectives, these fragmented evaluation protocols are underdeveloped to constitute a standard benchmark and conduct an inclusive evaluation. Recent efforts have attempted to systematically summarize and evaluate the characteristics of CRS, but they have mainly centered on establishing evaluation protocols through person-to-person conversation analysis and questionnaire interviews, lacking quantitative observations on existing CRS models Jin et al. [2023, 2021], Jannach [2022], Siro et al. [2023] and failing to systematically evaluate system-centric characteristics. Refer to Table 8 for better understanding our differences. In summary, existing protocols are considered too superficial and inadequate to uncover the shortcomings of CRS Jannach et al. [2021], Jannach [2023]. To this end, our Concept brings together system- and user-centric perspectives, enabling a more profound and inclusive evaluation protocol and ultimately contributing to the development of more practical CRS Li et al. [2023a].
@@
LLM as User Simulator. Manual evaluation often requires the user to constantly interact with CRS, which are labor-expensive and generally only feasible in an industry lab Huang et al. [2023], Gao et al. [2021]. To this end, LLM-based user simulators continue to be a rapidly evolving research area due to their labor-intensive nature and effectiveness Kocmi and Federmann [2023], Terragni et al. [2023], Tang et al. [2023], Cheng et al. [2023], Afzali et al. [2023]. A recent study on CRS also features interactive evaluation with LLM-based user simulation Wang et al. [2023d], experimentally demonstrating its effectiveness and reliability as a reliable alternative to humans. Taking inspiration from these studies, Concept also harnesses an LLM-based user simulator to interact with the CRS. Moreover, we equip our simulator with elaborately designed instructions and the Theory of Mind to mimic human social cognition Fischer [2023]. This encourages the simulator to reflect on its own predefined personality traits.
@@
LLM as Evaluator. A LLM-based evaluator could be seen as a combination of a LLM and its prompting strategy Zeng et al. [2023], Cohen et al. [2023], Chan et al. [2023], Wang et al. [2023b], Liu et al. [2023a]. Considering its flexibility, the instance-wise evaluator is widely adopted Ye et al. [2023], Wang et al. [2023c], Liu et al. [2023b]. Importantly, previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. [2023d]. These findings are further supported by recent work on CRS Wang et al. [2023d], resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. Inspired by them, we involve ability-specific scoring rubrics to gain reliable and aligned evaluations. We also encourage the evaluator to generate a rationale before assigning a score, further ensuring reliability Ye et al. [2023].
Appendix D Characteristics Categorization